The file oldstyle.dtx for use with L^{A}TEX 2ε.^{∗}

Robin Fairbairns

1995/05/04

1 Introduction

This file contains the external font information needed to load the cmmi and cmmib fonts for use to produce oldstyle numbers in L^{A}TEX text, together with a trivial L^{A}TEX 2ε package to enable them to be used in LaTeX text.

The package oldstyle implements two commands (one in L^{A}TEX 2.09 style — \oldstyle — and one in L^{A}TEX 2ε style — \textos).

The numbers look like this:

Command  Typeset text

               0123456789  0123456789

(normal numbers)

               \textos{0123456789}  0123456789
               \textbf{0123456789}  0123456789

(normal boldface)

               \textbf{%

                 \textos{0123456789}} 0123456789

I’ve not (yet) addressed the problems of using the oldstyle numbers in PostScript expert font sets; since there’s a lot of work going on in the area of PostScript use^{1}, I don’t believe that’s a suitable area for jiffy work such as this.

2 The .fd file

The .fd file specifies a version of the font family cmm in U encoding. This is the best I could think of (only the digits are oldstyle numbers, after all).

I have done what I believe to be the proper, and made oldstyle a new font shape, but I’m not (yet) sure about the ‘name’ (os) I’ve given it. Hence, this release is still only preliminary.

1 ⟨∗Ucmm⟩

 2 \DeclareFontFamily{U}{cmm}{\skewchar\font'177}
 3 \DeclareFontShape{U}{cmm}{m}{os}
 4  { <5> <6> <7> <8> <9> gen * cmmi
 5  <10><10.95>cmmi10
 6  <12><14.4><17.28><20.74><24.88>cmmi12

∗This file has version number v0.1, dated 1995/05/04.

1The betatest revised fontinst work, and Alan Jeffrey’s proposed update of fontinst itself

 7  }{}
 8 \DeclareFontShape{U}{cmm}{b}{os}{%
 9  <5> <6> <7> <8> <9> gen * cmmib
10  <10> <10.95> <12> <14.4> <17.28> <20.74> <24.88> cmmib10
11  }{}
12 \DeclareFontShape{U}{cmm}{bx}{os}
13 {<-> ssub * cmm/b/os}{}

14 ⟨/Ucmm⟩

3 The .sty file

The package is also trivial. It defines an oldstyle family and then declares commands to use it.

15 ⟨∗oldstyle⟩

16 \DeclareOldFontCommand {\oldstyle}{\usefont{U}{cmm}{m}{os}}%
17  {\mathos}
18 \DeclareTextFontCommand{\textos}{\oldstylefamily}
19 \DeclareMathAlphabet   {\mathos}{U}{cmm}{m}{os}

A series of constructs follows that enable the above to work; the constructs are mostly copied from the source of L^{A}TEX 2ε itself.

20 \DeclareRobustCommand\oldstylefamily{%
21  \not@math@alphabet\osshape\mathos
22  \usefont{U}{cmm}\f@series{os}}

23 ⟨/oldstyle⟩